OpenStudy (anonymous):

Choose the correct simplification of the expression. 4 / y to the -3 power.

OpenStudy (diyadiya):

\[\Large \frac{4}{y^{-3}} \ \ ??\]

OpenStudy (anonymous):

Yes

OpenStudy (diyadiya):

Note \[\huge x^{-m}= \frac{1}{x^m}\]

OpenStudy (anonymous):

I don't get that.

OpenStudy (diyadiya):

I'll give you an example

OpenStudy (anonymous):

Oh, so my equation/expression would actually be, 4 / -3?

OpenStudy (anonymous):

Oh, okay. (:

OpenStudy (diyadiya):

\[\Large x^{-2}= \frac{1}{x^2}\]

OpenStudy (diyadiya):

\[\Large x^{-6} = \frac{1}{x^6}\]

OpenStudy (diyadiya):

got that ?

OpenStudy (anonymous):

x = 1, which makes the 1. Then you take ou tht e - sign and re write as div. prob?

OpenStudy (diyadiya):

i didn't understand your question

OpenStudy (diyadiya):

brb

OpenStudy (diyadiya):

Back

OpenStudy (anonymous):

Hi. I understand your examples. a little.

OpenStudy (diyadiya):

when the variable is raised to negative power of a number ,it becomes 1/(variable)raised to positive power of that number

OpenStudy (anonymous):

I see.

OpenStudy (diyadiya):

Ok so \[\huge \frac{4}{y^{-3}}=\frac{4}{ \frac{1}{y^3}}= 4 \times \frac{y^3}{1}= 4y^3\] do you understand this ?

OpenStudy (anonymous):

All I NOTICE. Is that you took the negative out of the exponent.

OpenStudy (diyadiya):

when u take negative out of y^-3 ..it'll become 1/y^3 (same way as i did the examples)

OpenStudy (anonymous):

I see.

OpenStudy (diyadiya):

so got it ?

OpenStudy (anonymous):

\[4y^{2}\]

OpenStudy (anonymous):

not squared, cubed. \[4y^{3}\]

OpenStudy (diyadiya):

yeah its cubed

OpenStudy (anonymous):

So am right?

OpenStudy (anonymous):

The answer is, 4y^3?

OpenStudy (diyadiya):

Yes!

OpenStudy (anonymous):

I have one more.. it's more difficult though

OpenStudy (anonymous):

You want me to post it here or for everyone to see incase you don't understand?

OpenStudy (diyadiya):

post here its ok ..if i don't know i'll tell you

OpenStudy (anonymous):

Show the expanded form and simplification of \[x^{4} DIVIDED BY x^{2} Explain \in your own words how you can simplify.\]

OpenStudy (anonymous):

Explain in your own words how you can simplify.

OpenStudy (diyadiya):

\[\Large \frac{x^4}{x^2}\] Is that your question ?

OpenStudy (anonymous):

Yes. How do you get the division bar below x4?

OpenStudy (diyadiya):

\frac{numerator}{denominator}<<< in the equation box

OpenStudy (anonymous):

\[\frac{x^{4}}{x^{2}} \]

OpenStudy (diyadiya):

this is the same way how you did the previous problem .. forgot ?

OpenStudy (diyadiya):

\[\Large \frac{x^m}{x^n}=x^{m-n}\]

OpenStudy (anonymous):

\[\large\frac{x^{4}}{x^{2}}\] No, well yeah. I'm not that intelligen with remembering fractions nor am I with Algebra in general. :/ Sorry if I'm being a bother by being dumb. :/

OpenStudy (diyadiya):

Lol Don't Worry no problem at all :D

OpenStudy (anonymous):

Alright. :D

OpenStudy (diyadiya):

Remember this \[\Large \frac{x^m}{x^n}= x^{m-n}\]

OpenStudy (anonymous):

The answer would be \[x^{2}\] wouldn't it..?

OpenStudy (diyadiya):

Yes,You got it Right!! :) \[\Large x^2 \]

OpenStudy (anonymous):

:) only problem is that I have to explain it, in nothing but words. I can work it out, but she doesn't want that. She wants it written.

OpenStudy (anonymous):

Explain in your own words how you can simplify without having to write the expanded form.

OpenStudy (diyadiya):

Oh ,I don't think i can give you a *presentable* good explaination :( Sorry

OpenStudy (anonymous):

Doesn't have to be *presentable* just an explanation of how I would work it out with writing it in the expanded form.

OpenStudy (diyadiya):

let me see if i can get someone else to help you ..just a minute @lgbasallote Can you explain ?

OpenStudy (diyadiya):

@Mimi_x3 Can you :P >?

OpenStudy (anonymous):

I feel like an idiot. Honestly.

OpenStudy (diyadiya):

@Callisto

OpenStudy (anonymous):

Hello Callisto

OpenStudy (callisto):

Well... Just write by exponent law, x^m / x^n = x^(m-n) So, in this case, x^4 / x^2 = x^(4-2) = x^2

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

Wait. Part 2. .. :/ Create your own fraction with like bases, coefficients, and show its simplification.

OpenStudy (diyadiya):

@careless850 post it as new question this is lagging

OpenStudy (lgbasallote):

uhmm am i still needed? and yes this is laggy now :P

OpenStudy (anonymous):

I closed this. Look @ my new post for my other question.